Two ball are thrown simultaneously, A vertically upward with a speed of 20m/s from the ground , and B vertically downward from a height of 40m with the same speed along the same line of motion . At what point do two balls meet


1
Expert's answer
2021-12-07T09:48:56-0500

Let h be the distance from the ground where the collision takes place.

Distance from top of tower up to the collision point =40-h

Let collision took place after time t then:


For ball A

"h=20t-\frac{1}{2}gt^2……(1)"

For ball B

"40-h=20t+\frac{1}{2}gt^2……(2)"

Adding (1) and (2) we get

40=40t

Therefore t=1 sec

Putting value of t in (1) we get

"h=(20×1)-(\frac{1}{2}×9.8×1^2)=15.1m"

"( taking :g= 9.8 m \/s^2)"

h=15.1 metres

Therefore the two ball meet at 15.1m above the ground